OpenStudy (muddygirl):

A spinner has five sections, labeled A, B, C, D, and E. The spinner is spun 84 times, and the results are recorded in the table. What is the experimental probability of the spinner’s landing on A? Round to the nearest percent, if necessary. Outcome A B C D E Number of trials 14 20 18 15 17 A. 22% B. 20% C. 17% D. 14%.

OpenStudy (loginusernamed):

ok all of them added is 84

OpenStudy (loginusernamed):

84-17=Z Z-20=X X-15=Y Y-18=Anwser

OpenStudy (muddygirl):

Thank you!!